OpenStudy (anonymous):

Limits \[\lim_{x \rightarrow \infty } \sqrt{x-1} - \sqrt{x+2 }\] help plz

OpenStudy (bahrom7893):

limit as x approaches what?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty }\]

OpenStudy (bahrom7893):

\[\lim_{x \rightarrow + infinity}\] \[\sqrt{x-1} - \sqrt{x+2}\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty } \sqrt{x-1} - \sqrt{x+1 }\]

OpenStudy (bahrom7893):

Wolf works magic

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty } \sqrt{x-1} - \sqrt{x+2 }\]

OpenStudy (anonymous):

that is the question

OpenStudy (anonymous):

thanks my answer is correct